# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

### v1

@@ 3 One loop effective potential at zero temperature
We further define some finite counterterms to absorb terms which are independent of $\phi$ but otherwise finite,
$\delta\Lambda_{\rm fin.}=\frac{m^{4}}{64\pi^{2}}\left(\frac{1}{2}+\psi(2)\right),\qquad\delta\kappa_{\rm fin.}=\frac{m^{2}\psi(2)\left(\xi-\frac{1}{6}\right)}{32\pi^{2}}$
$\delta\alpha_{1,{\rm fin.}}=\frac{1}{5760\pi^{2}},\qquad\delta\alpha_{2,{\rm fin.}}=\frac{1}{5760\pi^{2}},\qquad\delta\alpha_{3,{\rm fin.}}=\frac{\left(\xi-\frac{1}{6}\right)^{2}}{4(4\pi)^{2}},$ (19)
to finally have the renormalised result
$V_{\rm{eff,Ren}}^{1-{\rm loop}}=\frac{1}{2}(m^{2}+\xi R)\phi^{2}+\frac{\lambda\phi^{4}}{4!}+\frac{\eta\phi^{3}}{3!}+\frac{1}{2(4\pi)^{2}}\left[\frac{m^{4}_{1}}{2}\left(\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{2}-\psi(2)\right)+\frac{m^{4}}{2}\left(\frac{1}{2}+\psi(2)\right)\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\left(\psi(2)(m^{2}_{1}-m^{2})-m_{1}^{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}\right)-\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\ln\frac{m^{2}_{1}}{4\pi\mu^{2}}\right]$ (20)
where $m_{1}^{2}=m^{2}+\frac{1}{2}\lambda\phi^{2}+\eta\phi$, and we have abbreviated
$f_{1}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-\frac{1}{3}a^{\lambda}\,_{\lambda}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}+\frac{1}{180}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-\frac{1}{180}R_{\mu\nu}R^{\mu\nu}$
@@
Let us now take the flat spacetime limit of 20, for a massless and quartic self interction. Taking $\psi(2)=0.422785$, we obtain the Coleman-Weinberg effective potential [5]
$V_{\rm{eff}}^{\rm flat}$ $=\frac{\lambda\phi^{4}}{4!}+\frac{\lambda^{2}\phi^{4}}{256\pi^{2}}\left[\ln\frac{\lambda\phi^{2}}{\mu^{2}}-\ln(8\pi)-\frac{1}{2}-0.422785\right]$ (21)
$=\frac{\lambda\phi^{4}}{4!}+\frac{\lambda^{2}\phi^{4}}{256\pi^{2}}\left[\ln\frac{\lambda\phi^{2}}{\mu^{2}}-4.146\right]$
We note a slight mismatch between the numerical factor appearing above ($4.146$), with that of [5], which is $4.166$.
@@ 4 The case at finite temparature
We next wish to generalise 12 for finite temperature, $T=\beta^{-1}$. In order to obtain the pure thermal contribution, we recall that one needs replacing the Euclidean momentum $k_{0}$ by discrete $\omega_{n}$ where $\omega_{n}=2\pi n/\beta$ for bosons, i.e. [20, 67]
@@
This leads to the replacements in 12
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]$
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{s}}\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\frac{1}{\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{s}}$ (23)
@@
$\Sigma=12\lambda\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{1}{\left(\vec{k}^{2}+m^{2}\right)^{\frac{1}{2}}}\frac{1}{e^{\beta\sqrt{\vec{k}^{2}+m^{2}}}-1}$ (24)
@@
$\Sigma=\lambda T^{2}\left(1-\frac{3m}{\pi T}\right)$ (25)
@@
${\bar{G}}(k)=\left[1-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\right]\frac{1}{k^{2}+m^{2}}$ (26)
@@
$\Sigma=\lambda T^{2}-\frac{3\lambda mT}{\pi}-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\left(\lambda T^{2}-\frac{3\lambda mT}{\pi}\right)=\lambda T^{2}-\frac{3\lambda mT}{\pi}+\left(\frac{1}{6}-\xi\right)\frac{3\lambda RT}{2\pi m}$ (27)
The third term of the final result exhibits mass generation in a thermal field theory in a weakly curved spacetime. We note that if the scalar is extremely light, the curvature correction term becomes very large, making the result untrustworthy. It seems that one needs to develop some non-perturbative resummation technique in order to tackle such issue. Nevertheless, the above result indeed indicates that the spacetime curvature can have interesting effects on thermal fields in a curved spacetime.
Let us now come to the computation of the effective potential. In 23, we may break the summation by separating the zero mode as
$\sum_{n=-\infty}^{\infty}=(n=0)+2\sum_{n=1}^{\infty}$ (28)
@@
Now, the pure thermal part of the first integral of 13 becomes at finite temperature
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]=\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}\Gamma\left(\frac{3-d}{2}\right)$ (29)
@@
$\sum_{n=1}^{\infty}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=\sum_{n=1}^{\infty}m_{1}^{d-1}\left(n\alpha\right)^{d-1}\left[1+\frac{1}{(n\alpha)^{2}}\right]^{\frac{d-1}{2}}$ (30)
@@
$m_{1}^{d-1}\alpha^{-\epsilon}\left[\alpha^{3}\zeta(-3)+\frac{3}{2}\alpha\zeta(-1)+\frac{3}{8}\alpha^{-1}\zeta(1+\epsilon)-\frac{1}{16}\alpha^{-3}\zeta(3)+\frac{1}{128}\alpha^{-5}\zeta(5)-\frac{5}{3072}\alpha^{-7}\zeta(3)+{\cal O}(\beta^{8})\right]$ (31)
where $\zeta(n)=\sum_{j=1}^{\infty}j^{-n}$ is the Riemann zeta function. Using the properties of this function pertaining to its regularisation, 29 becomes
@@
$+\frac{m^{6}_{1}\beta^{2}\zeta(3)}{768\pi^{4}}-\frac{m^{8}_{1}\beta^{4}\zeta(5)}{24576\pi^{6}}+\frac{m^{10}_{1}\beta^{6}\zeta(7)}{196608\pi^{8}}$ (32)
@@
$\left.+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{2048\pi^{6}}\zeta(5)+\frac{15m_{1}^{8}\beta^{6}}{24576\pi^{8}}\zeta(7)\right]$ (33)
@@
$\left.-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln{\frac{2\pi}{m_{1}\beta}}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{3m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{15m_{1}^{6}\beta^{6}}{3072\pi^{8}}\zeta(7)\right]$ (34)
An alternative and more rigorous derivation of the above integrals in the flat spacetime by breaking the Feynman propagator into temperature independent and dependent parts can be seen in [7]. We also note that in the three integrals appearing above there are a couple of temperature independent parts, such as the fifth term appearing on the right hand side of 32, even though 32, 33, 34 are pure thermal contributions. Such $\beta$ independent parts originate from the regularisation of the $\zeta$-function.
@@
$\left.+\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\left(\frac{1}{16\pi\beta m_{1}}-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln\frac{2\pi}{m_{1}\beta}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{5m_{1}^{6}\beta^{6}}{49152\pi^{8}}\zeta(7)\right)+{\cal O}(\beta^{7})\right]$
where $m_{1}^{2}=m^{2}+\lambda\phi^{2}/2+\eta\phi$, as earlier and $V_{\rm{eff}}^{1-{\rm loop}}(\beta^{-1}=0)$ is the zero temperature effective potential given by 20. Note that there is no temperature dependent divergence present in the above expression. This shows that renormalisation counterterms are independent of $\beta$, as is expected [25].
As we have mentioned earlier, the above one loop effective potential was first derived in [67] for $\eta=0$. However, our notation is different compared to this work. We also note from [67] that if we set the arbitrary mass scale $\mu=\beta^{-1}$ in 20, all the logarithmic terms appearing in the finite temperature part of the effective potential precisely cancel with that of the zero temperature. However, we wish to keep $\mu$ to be a free parameter in this work. Such terms will contribute non-trivially to the feature of the effective potential and hence to the spontaneous symmetry breaking phenomenon.
@@
$\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=m_{1}^{3}\left[1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}^{2}\beta}\right)^{4}+\cdots\right]$ (36)
Putting this expression back into 29, we obtain
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]_{\beta^{-1}\to 0}\approx\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2m_{1}^{3}}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left(1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}\beta}\right)^{4}\right)\Gamma\left(\frac{3-d}{2}\right)$
@@
$+\frac{3}{2}\left[-\frac{m_{1}^{3}}{3\pi\beta}+\left(\frac{1}{6}-\xi\right)\frac{m_{1}R}{2\pi\beta}+\left(-f_{1}+\frac{1}{4}\left(\frac{1}{6}-\xi\right)^{2}R^{2}\right)\frac{1}{\pi m_{1}\beta}\right]$ (38)
Setting $\beta^{-1}\to 0$ is possible in 38 whereas it is not possible in 35.
We further wish to consider two loop computations for the effective potential. However, before we do that, let us discuss the issue of spontaneous symmetry breaking and phase transition with the one loop high temperature effective potential, 35, we have found. We also note in 35 that for $\eta\neq 0$, and since both $\eta$ and $\phi$ can be positive or negative, the expression becomes invalid for $m_{1}^{2}\to 0$, owing to the $\ln m_{1}^{2}$ term. In other words, the effective potential will be valid in the parameter or the field space region where $m_{1}^{2}>0$. Owing to the fact that the background field $\phi$ is real, it is easy to see that this condition implies,
$\left(\frac{\eta}{m}\right)^{2}<2\lambda$ (39)

### v2

@@ 3 One loop effective potential at zero temperature
We further introduce counterterms to absorb terms which are independent of the background field $\phi$, but otherwise finite,
$\delta\Lambda_{\rm fin.}=\frac{m^{4}}{64\pi^{2}}\left(\psi(2)+\frac{1}{2}+\ln\frac{4\pi\mu^{2}}{m^{2}}\right),\qquad\delta\kappa_{\rm fin.}=\frac{m^{2}(\xi-1/6)}{32\pi^{2}}\left(\psi(2)+\ln\frac{4\pi\mu^{2}}{m^{2}}\right),$
$\delta\alpha_{1,{\rm fin.}}=\frac{1}{2880\pi^{2}}\left(\psi(1)+\ln\frac{4\pi\mu^{2}}{m^{2}}\right),\qquad\delta\alpha_{2,{\rm fin.}}=-\frac{1}{{2880\pi^{2}}}\left({\psi(1)}+\ln\frac{4\pi\mu^{2}}{m^{2}}\right),$
$\delta\alpha_{3,{\rm fin.}}=\frac{(\xi-1/6)^{2}}{{64\pi^{2}}}\left({\psi(1)}+\ln\frac{4\pi\mu^{2}}{m^{2}}\right).$ (20)
We will not absorb any field dependent finite terms by any counterterms, and will leave them intact. We thus finally have the renormalised result
$V_{\rm{eff,Ren}}^{1-{\rm loop}}=\frac{1}{2}(m^{2}+\xi R)\phi^{2}+\frac{\lambda\phi^{4}}{4!}+\frac{\eta\phi^{3}}{3!}+\frac{1}{2(4\pi)^{2}}\left[\frac{m^{4}_{1}}{2}\left(\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{2}-\psi(2)\right)+\frac{m^{4}}{2}\left(\frac{1}{2}+\psi(2)\right)-\frac{m^{4}}{2}\ln\frac{m^{2}}{4\pi\mu^{2}}\right.$
$\left.+\left(\frac{1}{6}-\xi\right)R\left(\psi(2)(m^{2}_{1}-m^{2})-m_{1}^{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}+m^{2}\ln\frac{m^{2}}{4\pi\mu^{2}}\right)-\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\ln\frac{m_{1}^{2}}{m^{2}}\right],$ (21)
where $m_{1}^{2}$ is given by 11, and we have abbreviated
$f_{1}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-\frac{1}{3}a^{\lambda}\,_{\lambda}=\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}+\frac{1}{180}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-\frac{1}{180}R_{\mu\nu}R^{\mu\nu}.$
Note that the renormalised result 21 does not contain any term independent of the background field $\phi$, and becomes vanishing at $\phi=0$. Thus we may regard it as a true effective potential for the scalar, containing the curvature effects.
@@
Let us now take the flat spacetime limit of 21, for a massless and quartic self interaction. Taking $\psi(2)\simeq 0.422785$ and setting $R=f_{1}=\eta=m=0$, we reproduce the Coleman-Weinberg effective potential [7]
$V_{\rm{eff}}^{\rm flat}=\frac{\lambda\phi^{4}}{4!}+\frac{\lambda^{2}\phi^{4}}{256\pi^{2}}\left[\ln\frac{\lambda\phi^{2}}{\mu^{2}}-4.146\right].$ (22)
Finally, we note that the derivation of the counterterms of 19 is completely equivalent to that of the imposition of renormalisation conditions. For example, for the mass renormalisation condition in curved spacetime we impose from 15, 18 and the tree level potential,
$\left(V_{\rm{eff,\ div.}}^{1-{\rm loop}}+\delta{\cal L}_{\rm matt}+\cal{L}_{\rm tree}\right)^{\prime\prime}_{\phi=0}=(m^{2}+\xi R)$
The above equation gives,
$\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\left[(m^{2}\lambda+\eta^{2})+\lambda\left(\xi-\frac{1}{6}\right)R\right]+(m^{2}+\delta m^{2})+(\xi+\delta\xi)R=m^{2}+\xi R,$
which immediately yields expressions same as that of 19 for $\delta m^{2}$ and $\delta\xi$. The other field counterterms can be derived in a similar manner. We further refer our reader to [57, 58] for a discussion on the equivalence of these two approaches in the context of $O(N)$ symmetric scalar field theory.
@@ 4 The case at finite temparature
We next wish to generalise 13 for finite temperature, $T=\beta^{-1}$. In order to obtain the pure thermal contribution, we recall that one needs replacing the Euclidean momentum $k_{0}$ by discrete $\omega_{n}$ where $\omega_{n}=2\pi n/\beta$ for bosons, i.e. [26, 74]
@@
This leads to the replacements in 13
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right],$
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{s}}\longrightarrow\sum_{n=-\infty}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\frac{1}{\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{s}}.$ (24)
@@
$\Sigma=12\lambda\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{1}{\left(\vec{k}^{2}+m^{2}\right)^{\frac{1}{2}}}\frac{1}{e^{\beta\sqrt{\vec{k}^{2}+m^{2}}}-1}.$ (25)
@@
$\Sigma=\lambda T^{2}\left(1-\frac{3m}{\pi T}\right).$ (26)
@@
${\bar{G}}(k)=\left[1-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\right]\frac{1}{k^{2}+m^{2}}.$ (27)
@@
$\Sigma=\lambda T^{2}-\frac{3\lambda mT}{\pi}-\left(\frac{1}{6}-\xi\right)R\frac{\partial}{\partial m^{2}}\left(\lambda T^{2}-\frac{3\lambda mT}{\pi}\right)=\lambda T^{2}-\frac{3\lambda mT}{\pi}+\left(\frac{1}{6}-\xi\right)\frac{3\lambda RT}{2\pi m}.$ (28)
The third term of the final result exhibits mass generation in a thermal field theory in a weakly curved spacetime. We note that if the scalar is extremely light, the curvature correction term becomes very large, making the result untrustworthy. It seems that one needs to develop some non-perturbative resummation technique in order to tackle such an issue. Nevertheless, the above result indeed indicates that the spacetime curvature can have interesting effects on thermal fields in a curved spacetime.
Let us now come to the computation of the effective potential. In 24, we may break the summation by separating the zero mode as
$\sum_{n=-\infty}^{\infty}\equiv(n=0)+2\sum_{n=1}^{\infty}.$ (29)
@@
Now, the pure thermal part of the first integral of 14 becomes at finite temperature
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]=\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}\Gamma\left(\frac{3-d}{2}\right).$ (30)
@@
$\sum_{n=1}^{\infty}\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=\sum_{n=1}^{\infty}m_{1}^{d-1}\left(n\alpha\right)^{d-1}\left[1+\frac{1}{(n\alpha)^{2}}\right]^{\frac{d-1}{2}},$ (31)
@@
$m_{1}^{d-1}\alpha^{-\epsilon}\left[\alpha^{3}\zeta(-3)+\frac{3}{2}\alpha\zeta(-1)+\frac{3}{8}\alpha^{-1}\zeta(1+\epsilon)-\frac{1}{16}\alpha^{-3}\zeta(3)+\frac{1}{128}\alpha^{-5}\zeta(5)-\frac{5}{3072}\alpha^{-7}\zeta(3)+{\cal O}(\beta^{8})\right],$ (32)
where $\zeta(n)=\sum_{j=1}^{\infty}j^{-n}$ is the Riemann zeta function. Using the properties of this function pertaining to its regularisation, 30 becomes
@@
$+\frac{m^{6}_{1}\beta^{2}\zeta(3)}{768\pi^{4}}-\frac{m^{8}_{1}\beta^{4}\zeta(5)}{24576\pi^{6}}+\frac{m^{10}_{1}\beta^{6}\zeta(7)}{196608\pi^{8}}.$ (33)
@@
$\left.+\frac{m_{1}^{4}\beta^{2}}{256\pi^{4}}\zeta(3)-\frac{m_{1}^{6}\beta^{4}}{6144\pi^{6}}\zeta(5)+\frac{5m_{1}^{8}\beta^{6}}{196608\pi^{8}}\zeta(7)\right]$ (34)
@@
$\left.-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln{\frac{2\pi}{m_{1}\beta}}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{5m_{1}^{6}\beta^{6}}{49152\pi^{8}}\zeta(7)\right].$ (35)
An alternative and more rigorous derivation of the above integrals in the flat spacetime by breaking the Feynman propagator into temperature independent and dependent parts can be seen in [9]. We also note that in the three integrals appearing above there are a couple of temperature independent parts, such as the fifth term appearing on the right hand side of 33, even though 33, 34, 35 are pure thermal contributions. Such $\beta$-independent parts originate from the regularisation of the $\zeta$-function.
@@
$\left.+\left(\frac{1}{2}\left(\frac{1}{6}-\xi\right)^{2}R^{2}-2f_{1}\right)\left(\frac{1}{16\pi\beta m_{1}}-\frac{1}{16\pi^{2}}\left(\psi(1)+\ln\frac{2\pi}{m_{1}\beta}\right)-\frac{m_{1}^{2}\beta^{2}}{128\pi^{4}}\zeta(3)+\frac{m_{1}^{4}\beta^{4}}{2048\pi^{6}}\zeta(5)-\frac{5m_{1}^{6}\beta^{6}}{49152\pi^{8}}\zeta(7)\right)+{\cal O}(\beta^{7})\right],$
where $m_{1}^{2}=m^{2}+\lambda\phi^{2}/2+\eta\phi$, as earlier and $V_{\rm{eff}}^{1-{\rm loop}}(\beta^{-1}=0)$ is the zero temperature effective potential given by 21. Note that there is no temperature dependent divergence present in the above expression. This shows that renormalisation counterterms are independent of $\beta$, as is expected [31]. Also, while investigating the behaviour of the above thermal effective potential, we will ignore all terms independent of $\phi$ but otherwise finite.
As we have mentioned earlier, the above one loop effective potential was first derived in [74] for $\eta=0$. However, our notation is different compared to this work. We also note from [74] that if we set the arbitrary mass scale $\mu=\beta^{-1}$ in 21, all the logarithmic terms appearing in the finite temperature part of the effective potential precisely cancel with that of the zero temperature. However, we wish to keep $\mu$ to be a free parameter in this work. Such terms will contribute non-trivially to the feature of the effective potential and hence to the spontaneous symmetry breaking phenomenon.
@@
$\left[\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]^{\frac{d-1}{2}}=m_{1}^{3}\left[1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}^{2}\beta}\right)^{4}+\cdots\right].$ (37)
Putting this expression back into 30, we obtain
$\sum_{n=1}^{\infty}\frac{1}{\beta}\int\frac{d^{d-1}\vec{k}}{(2\pi)^{d-1}}\ln\left[\vec{k}^{2}+\left(\frac{2\pi n}{\beta}\right)^{2}+m_{1}^{2}\right]_{\beta^{-1}\to 0}\approx\sum_{n=1}^{\infty}\frac{1}{\beta}\frac{2m_{1}^{3}}{(d-1)(4\pi)^{\frac{d-1}{2}}}\left(1+\frac{3}{2}\left(\frac{2\pi n}{m_{1}\beta}\right)^{2}+\frac{3}{8}\left(\frac{2\pi n}{m_{1}\beta}\right)^{4}\right)\Gamma\left(\frac{3-d}{2}\right).$
@@
$+\frac{3}{2}\left[-\frac{m_{1}^{3}}{3\pi\beta}+\left(\frac{1}{6}-\xi\right)\frac{m_{1}R}{2\pi\beta}+\left(-f_{1}+\frac{1}{4}\left(\frac{1}{6}-\xi\right)^{2}R^{2}\right)\frac{1}{\pi m_{1}\beta}\right].$ (39)
Setting $\beta^{-1}\to 0$ is possible in 39 whereas it is not possible in 36.
We further wish to consider two loop computations for the effective potential. However, before we do that, let us discuss the issue of spontaneous symmetry breaking and phase transition with the one loop high temperature effective potential, 36, we have found. We also note in 36 that for $\eta\neq 0$, and since both $\eta$ and $\phi$ can be positive or negative, the expression becomes invalid for $m_{1}^{2}\to 0$, owing to the $\ln m_{1}^{2}$ term. In other words, the effective potential will be valid in the parameter or the field space region where $m_{1}^{2}>0$. Owing to the fact that the background field $\phi$ is real, it is easy to see that this condition implies,
$\left(\frac{\eta}{m}\right)^{2}<2\lambda.$ (40)